THE DARK SINGLET STATE AS DOORWAY STATE OF INTERSYSTEM CROSSING IN DNA MONOMERS.

\underline{CHRISTIAN REICHARDT}, CARLOS E. CRESPO-HERNÁNDEZ, Case Western Reserve University, Department of Chemistry Center for Chemical Dynamics, 2080 Adelbert Road, Cleveland, Ohio 44106.

The excited state dynamics of 9-methylpurine (9MP) were studied with broadband transient absorption spectroscopy in the time regime from femtoseconds to 3 nanoseconds. Excitation of 9MP in aqueous solutions at 266 nm results in ultrafast internal conversion from the initially excited S_{2}(ππ^{∗}) state to the S_{1} state. Quantum chemical calculations that include bulk and explicit solvent interactions show that the S_{1} state has significant nπ^{∗} character. Population of the S_{1} state is followed by intersystem crossing (ISC) to the T_{1}(ππ^{∗}) state on a time scale of hundreds of picoseconds. Vanishingly small fluorescence yields were measured, supporting the dark character of the S_{1} state as well as the high triplet yield in 9MP. Analogous experiments in acetonitrile show a decrease in the ISC lifetime by almost 50 % but an equally high triplet yield. The results presented in this work demonstrate the important role that the dark singlet state has in modulating the excited-state dynamics of DNA monomers in solution.